A store mixes Brazillian coffee worth $10 per kilogram and Venezulen coffee worth $12 per kilogram. The mixture is to sell for $11 per kilogram. Find out how much of each should be used to make a 132-kilogram mixture?
Answer by stanbon(75887)

A store mixes Brazilian coffee worth $10 per kilogram and Venezulen coffee worth $12 per kilogram. The mixture is to sell for $11 per kilogram. Find out how much of each should be used to make a 132-kilogram mixture?
-----------------------------------
Equations:
Quantity Eq: b + v = 132 kg
Value Eq.:: 10b+12v = 11*132
------------------------------------
Multiply thru the 1st Eq by 10 to get:
10b + 10v = 10*132
-----
Subtract that from the Value Eq and solve for "v":
2v = 132
v = 66 kg (amt. of Venezualen coffee needed for the mixture)
----
Since b+v = 132, b = 132-66 = 66 kg (amt. of Brazilian coffee in the mix)
================================================
Cheers,
Stan H.